OpenStudy (anonymous):

Solve. m^-5=1/32 m =

OpenStudy (anonymous):

@flcon04

OpenStudy (anonymous):

I thinks it's 2

OpenStudy (flcon04):

ok give me a minute.

OpenStudy (anonymous):

k

OpenStudy (flcon04):

can you tell me how you found 2?

OpenStudy (anonymous):

yes

OpenStudy (flcon04):

okay tell me.

OpenStudy (anonymous):

D( m ) m = 0 m = 0 m = 0 m in (-oo:0) U (0:+oo) m^-5 = 1/32 // - 1/32 m^-5-(1/32) = 0 m^-5-1/32 = 0 1*m^-5 = 1/32 // : 1 m^-5 = 1/32 -5 < 0 1/(m^5) = 1/32 // * m^5 1 = 1/32*m^5 // : 1/32 32 = m^5 m^5 = 32 // ^ 1/5 m = 2 m = 2

OpenStudy (flcon04):

i dont think its 2 let me double check

OpenStudy (anonymous):

ok

OpenStudy (flcon04):

I double checked and its 2 you are right!

OpenStudy (anonymous):

Thank you! I have a couple more...

OpenStudy (flcon04):

ok i can do 2 more than i have to go

OpenStudy (anonymous):

What is the product written in scientific notation? (2.6 × 10–4)(2.2 × 108) A. 5.72 × 104 B. 5.72 × 10–32 C. 4.8 × 104 D. 4.8 × 10–2

OpenStudy (anonymous):

I think it's A

OpenStudy (flcon04):

Correct!

OpenStudy (anonymous):

Thank you... can I do 1 more

OpenStudy (flcon04):

yeah

OpenStudy (anonymous):

Solve. 3p/5=12 p =

OpenStudy (anonymous):

I think it's 20

OpenStudy (flcon04):

ok yeah you are right 3p/5=12 60/5 =12 so yeah its 20.

OpenStudy (flcon04):

if that helped a medal would be nice :)

OpenStudy (flcon04):

thanks!

OpenStudy (anonymous):

Thank you!!

OpenStudy (flcon04):

your welcome! i have to go now. but when I am on next time message me and ill help

OpenStudy (anonymous):

okay!!